OpenStudy (smoothini):

subtract and simplify (2p²q-3pq²+q³)-(p²q+q³)

OpenStudy (jjcoolja12):

Let's simplify step-by-step. 2p2q−3pq2+q3−(p2q+q3) Distribute the Negative Sign: =2p2q−3pq2+q3+−1(p2q+q3) =2p2q+−3pq2+q3+−1p2q+−1q3 =2p2q+−3pq2+q3+−p2q+−q3 Combine Like Terms: =2p2q+−3pq2+q3+−p2q+−q3 =(2p2q+−p2q)+(−3pq2)+(q3+−q3) =p2q+−3pq2 Answer: =p2q−3pq2 i dont know if this is actually right im a sponsor for mathpapa and im only able to use thier software